I neeed the dividened yield

Presented here are summarized data from the balance sheets and income statements of Wiper Inc.: WIPER INC. Condensed Balance Sheets December 31, 2020, 2019, 2018 (in millions) 2020 2019 2018 Current assets $ 683 $ 915 $ 763 Other assets 2,416 1,923 1,722 Total assets $ 3,099 $ 2,838 $ 2,485 Current liabilities $ 576 $ 806 $ 713 Long-term liabilities 1,513 1,003 851 Stockholders’ equity 1,010 1,029 921 Total liabilities and stockholders' equity $ 3,099 $ 2,838 $ 2,485 WIPER INC. Selected Income Statement and Other Data For the year Ended December 31, 2020 and 2019 (in millions) 2020 2019 Income statement data: Sales $ 3,053 $ 2,916 Operating income 299 313 Interest expense 87 68 Net income 200 195 Other data: Average number of common shares outstanding 41.6 47.0 Total dividends paid $ 53.0 $ 52.6 Required: Calculate return on investment, based on net income and average total assets, for 2020 and 2019. Calculate return on equity for 2020 and 2019. Calculate working capital and the current ratio for each of the past three years. Calculate earnings per share for 2020 and 2019. If Wiper's stock had a price/earnings ratio of 13 at the end of 2020, what was the market price of the stock? Calculate the cash dividend per share for 2020 and the dividend yield based on the market price calculated in part e. Calculate the dividend payout ratio for 2020. Assume that accounts receivable at December 31, 2020, totaled $312 million. Calculate the number of days' sales in receivables at that date. Calculate Wiper's debt ratio and debt/equity ratio at December 31, 2020 and 2019. Calculate the times interest earned ratio for 2020 and 2019.

I neeed the dividened yield
Answer #1

Cash dividends per share = total cash dividends / weighted average shares outstanding = 53/41.6

= $1.27

Dividend yield = cash dividends per share / market price per share

2020 Earnings per share = net income/ weighted average number of shares outstanding = 200/41.6 = $4.81

2020 market price =price/earnings ratio * earnings per share = 13*4.81 = $62.53

Therefore,

Dividend yield = 1.27/62.53 = 2%